### v1

@@ Abstract
We focus on narrowing this gap by revisiting and optimizing DeepImpact, a sparse retrieval approach that uses DocT5Query for document expansion followed by a BERT language model to learn impact scores for document terms. We first reinvestigate the expansion process and find that the recently proposed Doc2Query-- query filtration does not enhance retrieval quality when used with DeepImpact. Instead, substituting T5 with a fine-tuned Llama 2 model for query prediction results in a considerable improvement. Subsequently, we study training strategies that have proven effective for other models, in particular the use of hard negatives, distillation, and pre-trained CoCondenser model initialization. Our results significantly narrow the effectiveness gap with the most effective versions of SPLADE.
@@ 1 Introduction
Early in the development of sparse methods, Dai and Callan [7] introduced DeepCT, which leverages contextual word representations from BERT to generate more effective document term weights that can be stored and used in inverted indexes. DeepCT estimates a term’s context-specific score by learning to map its final BERT hidden state to a single term weight. However, one of the main limitations of DeepCT is that it still suffers from the vocabulary mismatch problem, as it only assigns weights to terms present in the document. Furthermore, as highlighted by Mallia et al. [8], DeepCT is trained as a per-token regression task, which requires knowing the ground truth for each term weight, and thereby prevents individual scores to co-adapt for identifying relevant documents.
@@
In another important line of work, Formal et al. [12] introduced SPLADE, a sparse retrieval model that improves upon the key ideas of SparTerm by adding a log-saturation effect and changes to the training mechanism. It predicts the term importance for all the tokens in the BERT WordPiece Tokenizer vocabulary. SPLADE has been continuously optimized over the last few years, most recently with the release of SPLADE-v3 [13], which obtains significant improvements in quality over earlier versions. Other important approaches for sparse learned indexing include COIL [14], the uniCOIL variant by Lin and Ma [15], and TILDE [16].
@@ 1.1 Motivation and Our Contributions
Since the publication of DeepImpact [8], a number of techniques have been introduced to advance learned sparse retrieval. These improvements include the use of more powerful large language models (LLMs), the idea of filtering queries generated for document expansion [17], enhanced training methodologies such as the CoCondenser approach [18], the use of hard negatives [19, 20, 21], and distillation for transferring knowledge learned from powerful cross-encoder models into simpler and more efficient models [22, 23, 24, 25]. The more recent versions of SPLADE [22, 13] have already adopted some of these advances, resulting in impressive performance gains.
Our objective in this paper is to incorporate these cutting-edge techniques into the DeepImpact framework and evaluate the resulting performance gains and tradeoffs. The enhanced version, which we call DeeperImpact  achieves substantial improvements in retrieval metrics over the earlier version [8], and significantly closes the performance gap with the recent versions of SPLADE. This is achieved by using Llama 2 for document expansion, initializing with the CoCondenser model, and incorporating hard negatives and distillation during training.
@@ 2 DeeperImpact
In this section, we provide a detailed explanation of the modifications we attempt in the document expansion strategy and training approach of the original DeepImpact model to enhance its performance.
@@ 2.1.1 Doc2Query--
We investigate whether the hallucinations in DocT5Query affect DeepImpact’s performance. To integrate the Doc2Query-- filtration mechanism into the DeepImpact framework, we first retrieve 80 prospective expansion queries for each document generated using the DocT5Query method  [^1] https://huggingface.co/datasets/macavaney/d2q-msmarco-passage. These candidate queries are then evaluated using an ELECTRA-based relevance model—mapping each query-document pair to a real-valued relevance score. We utilize the pre-computed relevance scores published by the authors [^2] https://huggingface.co/datasets/macavaney/d2q-msmarco-passage-scores-electra. Specifically, we retain only those queries that achieve a relevance score within the top 30 percentile globally. This selective filtration process ensures that only the most relevant and contextually accurate queries are included in the final expanded document set.
This experiment essentially serves as an evaluation of DeepImpact’s robustness. DeepImpact—as a sparse retrieval model that learns to assign impact scores to each term in a document—inherently functions as a relevance model. This implies that, ideally, DeepImpact should be capable of suppressing the scores of non-relevant expansion terms introduced as a result of hallucinations from the DocT5Query expansion process. Consequently, the integration of the Doc2Query-- filtration mechanism should not lead to significant performance improvements if DeepImpact has been trained correctly. However, should such improvements be observed, this may indicate potential shortcomings in the initial training process or highlight opportunities for further refinement of the model.
@@ 2.1.2 Llama 2 Expansions
We use the same dataset as DocT5Query for fine-tuning the pre-trained Llama 2 model i.e. 532k document-query pairs from MSMARCO Passage Qrels Train Dataset [28]. During inference, we use a combined Top-K and Top-P sampling technique for generating the queries—first selecting the $K$ most likely tokens from the model’s vocabulary, then reducing it to the smallest set whose cumulative probability exceeds the threshold $P$, and finally sampling from this reduced set. This balances guiding the model towards probable outcomes while maintaining query diversity
@@
The overarching objective of modifying the expansion model is to generate more relevant and contextually accurate terms during the document expansion phase. DeepImpact then subsequently learns to score the terms that were not present in the original document. By increasing the quality and relevance of these queries for document expansion, we aim to reduce the vocabulary mismatch problem further, and thereby achieve an improvement in retrieval performance over the original DeepImpact model.
@@ 2.2.1 Pre-Trained CoCondenser Checkpoint
Recent advancements in the field of Natural Language Processing (NLP) have incorporated the principles of contrastive learning in order to generate high-quality document representations without the need of manual annotations [18, 29, 30, 31]. A prominent approach in Information Retrieval that illustrates this trend is the CoCondenser model [18], which leverages the Condenser pre-training architecture [32] in conjunction with a contrastive loss function. This dual approach facilitates the learning of a more refined embedding space from different spans of documents.
@@ 2.2.2 Hard Negatives
One of the primary concerns in training retrieval models is the quality of the negative samples included in the training data. Traditionally, the training of neural rankers, including the original DeepImpact model, has relied heavily on the MSMARCO Triples dataset, which comprises triples of the form $(q,~{}d^{+},~{}d^{-})$, where $q$ is a query, $d^{+}$ is a relevant (positive) document, and $d^{-}$ is a non-relevant (negative) document. These negatives are typically sampled using BM25 scores. While BM25 is a robust lexical matching algorithm, it does not fully capture the semantic nuances between relevant and non-relevant documents that are syntactically very similar. Thus, the negatives identified by BM25 often lack the semantic richness that more advanced dense retrieval models are capable of identifying. This limitation can lead to sub-optimal training and subsequently poorer retrieval performance.
We first experimented with the common trick of using multiple in-batch negatives instead of triples. Now, every query $q_{i}$ in a batch of triples of size $|B|$ will have one positive document $d_{i}^{+}$, one negative document $d_{i}^{-}$, and a set of $|B|-1$ additional negative documents $\{d_{j}^{+}\}_{j\neq i}$ that are positive documents for other queries in the batch. However, we observed only negligible improvements in the performance of DeepImpact using in-batch negatives.
@@ 2.2.3 Distillation
Distillation has become a popular method for enhancing the performance of neural rankers in information retrieval tasks. This technique has been extensively validated across both sparse as well as dense retrieval systems [22, 23, 24, 25]. Although computationally expensive for large-scale retrieval tasks, cross-encoder models, by design, capture more intricate interactions between the query and document, which makes them ideal teacher models for distillation.
In addition to using hard negatives, we integrate knowledge distillation from a cross-encoder teacher model to further enhance DeepImpact’s performance. Again, we make use of the msmarco-hard-negatives dataset which provides not only the Top-50 hard negatives and positives for each training query, but also the relevance scores associated with each pair. These scores are derived from the ms-marco-MiniLM-L-6-v2 cross-encoder model [^4] https://huggingface.co/cross-encoder/ms-marco-MiniLM-L-6-v2.
@@
As an alternative, we also experimented with Margin-MSE loss for distillation, where we optimize the mean squared error of the margin between relevant and non-relevant documents as predicted by the teacher and student models. This approach focuses on relative score differences rather than absolute values. However, our empirical results demonstrate that KL divergence loss yields better performance, leading us to adopt it as the primary distillation method.
@@ 3 Experimental Results
Baselines. In our experiments, we compare our proposed DeeperImpact model against uniCOIL with TILDE document expansion, SPLADE (which refers to CoCondenser-EnsembleDistil in [22]), and the original DeepImpact model.
@@
Overall comparison. Our first experiment aims to compare the retrieval effectiveness of the proposed DeeperImpact model with the original DeepImpact, uniCOIL with TILDE document expansion, SPLADE (which refers to CoCondenser-EnsembleDistil in [22]), and SPLADE-v3. We also report the retrieval effectiveness of DeepImpact when combined with Doc2Query--, which we will refer to as DeepImpact--. The results are presented in Table 1, which shows effectiveness and efficiency for the three query logs on MSMARCO. We retrieve the top 1000 documents for each query, without re-ranking, and report the values of NDCG@10, MRR@10, and Recall@1k, as well as MRT.
@@
We first notice that integrating Doc2Query-- filtration into the DeepImpact framework does not substantially improve its performance, but rather hampers it in most of the cases. This strengthens our claim that DeepImpact is already trained correctly as a relevance/filtration model. Instead of filtering out the predicted queries, it makes more sense to provide DeepImpact with a larger set of expanded documents regardless of the fact that it might contain terms predicted as a result of hallucination.
@@
We also notice that DeeperImpact demonstrates significant improvements on the MSMARCO Dev Queries, TREC 2019, and TREC 2020 queries compared to the original DeepImpact model. Specifically for the MSMARCO Dev Queries, DeeperImpact achieves an NDCG@10 score of 43.49, which is notably higher than the 38.65 of DeepImpact. Additionally, DeeperImpact shows better performance in MRR@10 (37.30 versus 32.74) and Recall@1k (96.83 versus 94.76). These improvements can be attributed to the better expansion strategy with Llama 2, incorporation of hard negatives and distillation during training, and the use of the CoCondenser checkpoint—which we will look at in detail in the subsequent sections.
@@
SPLADE consistently performs better in terms of the MRR@10, Recall@1k, and NDCG@10 metrics compared to the DeeperImpact model. However, while SPLADE remains dominant in terms of retrieval effectiveness, it is significantly less efficient than DeeperImpact, with an MRT of 378.00 ms compared to 39.62 ms for DeeperImpact on the MSMARCO Dev Queries. This trend is also consistent across the TREC 2019 and TREC 2020 datasets.
@@
These results underscore the retrieval effectiveness of the DeeperImpact model while continuing to maintain high retrieval efficiency. This balance is particularly advantageous for large-scale information retrieval tasks, where computational efficiency and response time are critical, making DeeperImpact a superior alternative to models like SPLADE.
@@
Table 1: Effectiveness metrics and mean response time (MRT, in ms) for first-stage methods, on MSMARCO Dev Queries, TREC 2019 queries, and TREC 2020 queries.
@@
Ablation study. We also conduct an ablation study to rigorously evaluate the contributions of various components in enhancing the performance of our proposed DeeperImpact model. We begin with a simplified version of DeepImpact without any document expansions and trained using the MSMARCO Triples, referred to as DeeperImpactBase. Subsequently, we augment this base model with the following series of improvements:
@@
    + Llama 2: Incorporates document expansions generated by the fine-tuned Llama 2 model.
@@
The results are summarized in Table 2. Our performance metrics include MRR@10, Recall@1k, and MRT for the MSMARCO Dev Queries. We notice incremental improvements in the MRR@10 as we continue adding the refinements. The most significant improvement in MRR@10 and Recall@1k is observed when incorporating LLama2 expansions, which can be attributed to its effectiveness in addressing vocabulary mismatch issues.
Interestingly, subsequent enhancements with CoCondenser and distillation do not further improve recall, suggesting that even LLama2 expansions trained with the triples are as effective in capturing the relevant documents. However, these enhancements do contribute to better MRR@10, indicating improved ranking precision. This implies that while document expansion significantly aids in ensuring a broader recall of relevant documents, the fine-tuning with CoCondenser and distillation techniques improves the ranking, likely due to better contextualized representations and more robust handling of difficult queries.
@@
Table 2: Ablation study on MSMARCO Dev Queries.
@@
DocT5Query vs. Llama 2 Expansions In this section, we compare the quality of document expansions generated by DocT5Query and Llama 2 using BM25 scores on the MSMARCO Dev Queries dataset. The effectiveness metrics are summarized in Table 3 for varying numbers of queries appended to each document in the corpus.
@@
Table 3: Effectiveness metrics on MSMARCO Dev Queries using BM25 scores after DocT5Query vs Llama 2 expansions for n queries
MRR@10
DocT5Query Llama 2
n = 5 25.9 25.4
n = 10 26.5 26.5
n = 20 27.2 25.4
n = 40 27.7 27.4
n = 80 27.8 27.4
Recall@1000
DocT5Query Llama 2
n = 5 92.9 94.0
n = 10 93.9 94.8
n = 20 94.4 95.4
n = 40 94.7 95.5
n = 80 94.5 95.6
Recall at different depths
Table 4 presents the recall metrics at various depths on the MSMARCO Dev Queries dataset for four different models: uniCOIL (with TILDE document expansion), SPLADE (which refers to CoCondenser-EnsembleDistil in [22]), DeepImpact  and DeeperImpact. We notice that SPLADE consistently achieves the highest recall at all the depths. However, DeeperImpact—which offers much faster query processing—also demonstrates strong performance, keeping ahead of uniCOIL in all the depths.
Remarkably, DeeperImpact attains a recall at depth 400 that is equivalent to the Recall@1000 achieved by the original DeepImpact model. This highlights the efficiency of DeeperImpact in retrieving relevant documents at shallower depths, demonstrating its effectiveness in reducing the number of retrieved documents necessary to achieve high recall.
Table 4: Recall at different depths on MSMARCO Dev Queries

### v2

@@ Abstract
We focus on narrowing this gap by revisiting and optimizing DeepImpact, a sparse retrieval approach that uses DocT5Query for document expansion followed by a BERT language model to learn impact scores for document terms. We first reinvestigate the expansion process and find that the recently proposed Doc2Query-- query filtration does not enhance retrieval quality when used with DeepImpact. Instead, substituting T5 with a fine-tuned Llama 2 model for query prediction results in a considerable improvement. Subsequently, we study training strategies that have proven effective for other models, in particular the use of hard negatives, distillation, and pre-trained CoCondenser model initialization. Our results substantially narrow the effectiveness gap with the most effective versions of SPLADE.
@@ 1 Introduction
1.1 Related Works
Early in the development of sparse methods, Dai and Callan [7] introduced DeepCT, which uses contextual word representations from BERT to generate more effective document term weights that can be stored and used in inverted indexes. DeepCT estimates a term’s context-specific score by learning to map its final BERT hidden state to a single term weight. However, one of the main limitations of DeepCT is that it still suffers from the vocabulary mismatch problem, as it only assigns weights to terms present in the document. Furthermore, as highlighted by Mallia et al. [8], DeepCT is trained as a per-token regression task, which requires knowing the ground truth for each term weight, and thereby prevents individual scores to co-adapt for identifying relevant documents.
@@ 1.1 Related Works
In another important line of work, Formal et al. [12] introduced SPLADE, a sparse retrieval model that improves upon the key ideas of SparTerm by adding a log-saturation effect and changes to the training mechanism. It predicts the term importance for all the tokens in the BERT WordPiece Tokenizer vocabulary. SPLADE has been continuously optimized over the last few years, most recently with the release of SPLADE-v3 [13], which obtains significant improvements in quality over earlier versions.
Other important approaches for sparse learned indexing include COIL [14], the uniCOIL variant by Lin and Ma [15], and TILDE [16, 17]. COIL [14] uses a deep language model to encode query and document tokens into contextualized vectors, stored directly in the inverted index. The retrieval process involves calculating inner products between matching token vectors, combining the efficiency of exact lexical match with the enhanced representational capabilities of LLMs. uniCOIL [15] is a simplified variant of COIL that reduces token dimensions to scalar weights, making it directly comparable to DeepCT and DeepImpact. When combined with document expansion techniques like DocT5Query, uniCOIL achieves effectiveness comparable to COIL.
Meanwhile, TILDE [16] follows a query likelihood paradigm, estimating relevance of a passage to a query based on the likelihood of generating the query text from the passage. However, this results in large index sizes as it requires storing inference on the entire BERT vocabulary. TILDEv2 [17] improves upon this by abandoning query likelihood estimation and employing exact term matching. It uses BERT to output a scalar importance weight for tokens only in the passage. In addition, TILDEv2 performs passage expansion using the original TILDE’s query token likelihood distribution, significantly speeding up the expansion process compared to auto-regressive query prediction methods like docT5query.
@@ 1.2 Motivation and Our Contributions
Since the publication of DeepImpact [8], a number of techniques have been introduced to advance learned sparse retrieval. These improvements include the use of more powerful large language models (LLMs), the idea of filtering queries generated for document expansion [18], enhanced training methodologies such as the CoCondenser approach [19], the use of hard negatives [20, 21, 22], and distillation for transferring knowledge learned from powerful cross-encoder models into simpler and more efficient models [23, 24, 25, 26]. The more recent versions of SPLADE [23, 13] have already adopted some of these advances, resulting in impressive effectiveness gains.
In this paper, our objective is to incorporate these cutting-edge techniques into the DeepImpact framework and evaluate the resulting gains in effectiveness and efficiency, along with the associated trade-offs. The enhanced version, which we call DeeperImpact  achieves significant improvements in retrieval metrics over the earlier version [8], and substantially closes the effectiveness gap with the recent versions of SPLADE. This is achieved by using Llama 2 for document expansion, initializing with the CoCondenser model, and incorporating hard negatives and distillation during training.
@@ 2 DeeperImpact
In this section, we provide a detailed explanation of the modifications we attempt in the document expansion strategy and training approach of the original DeepImpact model to enhance its effectiveness.
@@ 2.1 Document Expansion
This subsection explores methods to improve the document expansion phase, which is crucial for addressing the vocabulary mismatch problem and enhancing the overall retrieval effectiveness. Specifically, we address the challenges associated with hallucinated content in the DocT5Query method and consider the adoption of a more advanced language model for more relevant and contextually accurate query generation.
@@ 2.1.1 Doc2Query--
We investigate whether the hallucinations in DocT5Query affect DeepImpact’s effectiveness and efficiency. To integrate the Doc2Query-- filtration mechanism into the DeepImpact framework, we first retrieve 80 prospective expansion queries for each document generated using the DocT5Query method. [^1] https://huggingface.co/datasets/macavaney/d2q-msmarco-passage These candidate queries are then evaluated using an ELECTRA-based relevance model—mapping each query-document pair to a real-valued relevance score. We utilize the pre-computed relevance scores published by the authors. [^2] https://huggingface.co/datasets/macavaney/d2q-msmarco-passage-scores-electra Specifically, we retain only those queries that achieve a relevance score within the top 30 percentile of all expansion queries across the entire dataset. The choice of top 30% threshold was based on the findings of the original paper. This selective filtration process ensures that only the most relevant and contextually accurate queries are included in the final expanded document set.
This experiment essentially serves as an evaluation of DeepImpact’s robustness. DeepImpact—as a sparse retrieval model that learns to assign impact scores to each term in a document—inherently functions as a relevance model. This implies that, ideally, DeepImpact should be capable of suppressing the scores of non-relevant expansion terms introduced as a result of hallucinations from the DocT5Query expansion process. Consequently, the integration of the Doc2Query-- filtration mechanism should not lead to substantial improvements in the effectiveness if DeepImpact has been trained correctly. However, should such improvements be observed, this may indicate potential shortcomings in the initial training process or highlight opportunities for further refinement of the model.
@@ 2.1.2 Llama 2 Expansions
We use the same dataset as DocT5Query for fine-tuning the pre-trained Llama 2 model i.e. 532k document-query pairs from MS MARCO Passage Qrels Train Dataset [29]. During inference, we use a combined top-k and top-p sampling technique for generating the queries—first selecting the $k$ most likely tokens from the model’s vocabulary, then reducing it to the smallest set whose cumulative probability exceeds the threshold $p$, and finally sampling from this reduced set. This balances guiding the model towards probable outcomes while maintaining query diversity.
@@
The overarching objective of modifying the expansion model is to generate more relevant and contextually accurate terms during the document expansion phase. DeepImpact then subsequently learns to score the terms that were not present in the original document. By increasing the quality and relevance of these queries for document expansion, we aim to reduce the vocabulary mismatch problem further, and thereby achieve an improvement in retrieval effectiveness over the original DeepImpact model.
@@ 2.2 Training Improvements
We have made several key improvements to enhance our model’s training, including the use of a pre-trained CoCondenser checkpoint, the incorporation of hard negatives, and the application of knowledge distillation. Below, we provide a detailed discussion of these improvements.
@@ 2.2.1 Pre-Trained CoCondenser Checkpoint
Recent advancements in the field of Natural Language Processing (NLP) have incorporated the principles of contrastive learning in order to generate high-quality document representations without the need of manual annotations [19, 30, 31, 32]. A prominent approach in Information Retrieval that illustrates this trend is the CoCondenser model [19], which uses the Condenser pre-training architecture [33] in conjunction with a contrastive loss function. This dual approach facilitates the learning of a more refined embedding space from different spans of documents.
@@ 2.2.2 Hard Negatives
One of the primary concerns in training retrieval models is the quality of the negative samples included in the training data. Traditionally, the training of neural rankers, including the original DeepImpact model, has relied heavily on the MS MARCO Triples dataset, which comprises triples of the form $(q,~{}d^{+},~{}d^{-})$, where $q$ is a query, $d^{+}$ is a relevant (positive) document, and $d^{-}$ is a non-relevant (negative) document. These negatives are typically sampled using BM25 scores. While BM25 is a robust lexical matching algorithm, it does not fully capture the semantic nuances between relevant and non-relevant documents that are syntactically very similar. Thus, the negatives identified by BM25 often lack the semantic richness that more advanced dense retrieval models are capable of identifying. This limitation can lead to sub-optimal training and subsequently poorer retrieval effectiveness.
We first experimented with the common trick of using multiple in-batch negatives instead of triples. Now, every query $q_{i}$ in a batch of triples of size $|B|$ will have one positive document $d_{i}^{+}$, one negative document $d_{i}^{-}$, and a set of $|B|-1$ additional negative documents $\{d_{j}^{+}\}_{j\neq i}$ that are positive documents for other queries in the batch. However, we observed only negligible improvements in the effectiveness of DeepImpact using in-batch negatives.
@@ 2.2.3 Distillation
Distillation, a technique for transferring knowledge from a powerful teacher model to a simpler student model, has become a popular method for enhancing the effectiveness of neural rankers in information retrieval tasks. This technique has been extensively validated across both sparse as well as dense retrieval systems [23, 24, 25, 26]. Although computationally expensive for large-scale retrieval tasks, cross-encoder models, by design, capture more intricate interactions between the query and document, which makes them ideal teacher models for distillation.
In addition to using hard negatives, we integrate knowledge distillation from a cross-encoder teacher model to further enhance DeepImpact’s performance. Again, we make use of the msmarco-hard-negatives dataset which provides not only the Top-50 hard negatives and positives for each training query, but also the relevance scores associated with each pair. These scores are derived from the ms-marco-MiniLM-L-6-v2 cross-encoder model. [^4] https://huggingface.co/cross-encoder/ms-marco-MiniLM-L-6-v2
@@
As an alternative, we also experimented with Margin-MSE loss for distillation, where we optimize the mean squared error of the margin between relevant and non-relevant documents as predicted by the teacher and student models. This approach focuses on relative score differences rather than absolute values. However, our empirical results demonstrate that KL divergence loss yields better performance, leading us to adopt KL divergence loss as the primary distillation method.
@@ 3 Experimental Results
Baselines. In our experiments, we compare our proposed DeeperImpact model against uniCOIL with TILDE document expansion, SPLADE (which refers to CoCondenser-EnsembleDistil in [23]), the recent SPLADE-v3 [13] and the original DeepImpact model.
@@ 3.1 Overall comparison
Our first experiment aims to compare the retrieval effectiveness of the proposed DeeperImpact model with the original DeepImpact, uniCOIL with TILDE document expansion, SPLADE (which refers to CoCondenser-EnsembleDistil in [23]), and SPLADE-v3. We also report the retrieval effectiveness of DeepImpact when combined with Doc2Query--, which we will refer to as DeepImpact--. The results are presented in Table 1, which shows effectiveness and efficiency for the three query logs on MS MARCO. We retrieve the top 1000 documents for each query, without re-ranking, and report the values of NDCG@10, MRR@10, and Recall@1k, as well as MRT.
@@
We first notice that integrating Doc2Query-- filtration into the DeepImpact framework does not substantially improve its effectiveness and in some cases results in marginal declines. This strengthens our claim that DeepImpact is already trained correctly as a relevance/filtration model. Instead of filtering out the predicted queries, it makes more sense to provide DeepImpact with a larger set of expanded documents regardless of the fact that they might contain terms predicted as a result of hallucination.
@@
We also notice that DeeperImpact demonstrates substantial improvements on the MS MARCO Dev Queries, TREC 2019, and TREC 2020 queries compared to the original DeepImpact model. Specifically for the MS MARCO Dev Queries, DeeperImpact achieves an NDCG@10 score of 43.49, which is notably higher than the 38.65 of DeepImpact. Additionally, DeeperImpact shows better effectiveness in MRR@10 (37.30 versus 32.74) and Recall@1k (96.83 versus 94.76). These improvements can be attributed to the better expansion strategy with Llama 2, incorporation of hard negatives and distillation during training, and the use of the CoCondenser checkpoint—which we will look at in detail in the subsequent sections.
It is worth noting that while uniCOIL performs worse both in terms of effectiveness and efficiency on the MS MARCO Dev Queries, it achieves better effectiveness on the TREC 2019 and TREC 2020 queries, surpassing DeeperImpact in MRR@10 and Recall@1k. Additionally, it has a much lower MRT for both TREC 2019 and TREC 2020 queries. This suggests that uniCOIL may be more effective for certain types of queries within these datasets.
@@
SPLADE consistently performs better in terms of the MRR@10, Recall@1k, and NDCG@10 metrics compared to the DeeperImpact model. However, while SPLADE remains dominant in terms of retrieval effectiveness, it is substantially less efficient than DeeperImpact, with an MRT of 378.00 ms compared to 39.62 ms for DeeperImpact on the MS MARCO Dev Queries. This trend is also consistent across the TREC 2019 and TREC 2020 queries.
@@
These results underscore the retrieval effectiveness of the DeeperImpact model while continuing to maintain high retrieval efficiency. This balance is particularly advantageous for large-scale information retrieval tasks, where computational efficiency and response time are critical. Given these strengths, DeeperImpact presents a compelling alternative to models like SPLADE for scenarios where both effectiveness and efficiency are important. However, if the primary focus is solely on retrieval effectiveness, SPLADE may still be preferred.
@@
Table 1: Effectiveness metrics and mean response time (MRT, in ms) for first-stage methods, on MS MARCO Dev Queries, TREC 2019 queries, and TREC 2020 queries. For the recent SPLADE-v3, only the scores available in their paper have been reported.
@@
3.2 Ablation Study
We also conduct an ablation study to rigorously evaluate the contributions of various components in enhancing the effectiveness of our proposed DeeperImpact model. We begin with a simplified version of DeepImpact without any document expansions and trained using the MS MARCO Triples, referred to as DeeperImpactBase. Subsequently, we augment this base model with the following series of improvements:
@@ 3.2 Ablation Study
    + Llama 2 expansions: Incorporates document expansions generated by the fine-tuned Llama 2 model.
@@
The results are summarized in Table 2. Our performance metrics include MRR@10, Recall@1k, and MRT for the MS MARCO Dev Queries. We notice incremental improvements in the MRR@10 as we continue adding the refinements. The most substantial improvement in MRR@10 and Recall@1k is observed when incorporating Llama2 expansions, which can be attributed to its effectiveness in addressing vocabulary mismatch issues. We can further compare these metrics to those reported using other models like DocT5Query (DeepImpact) and Doc2Query-- (DeepImpact--) in Table 1, and observe that using Llama2 expansions demonstrate better retrieval effectiveness in both metrics.
Interestingly, subsequent enhancements with CoCondenser and distillation do not further improve recall, suggesting that even Llama2 expansions trained with the triples are as effective in capturing the relevant documents. However, these enhancements do contribute to better MRR@10, indicating improved ranking precision. This implies that while document expansion substantially aids in ensuring a broader recall of relevant documents, the fine-tuning with CoCondenser and distillation techniques improves the ranking, likely due to better contextualized representations and more robust handling of difficult queries.
@@
Table 2: Ablation study on MS MARCO Dev Queries evaluating effectiveness metrics and mean response time (MRT, in ms) as components are incrementally added to DeepImpact.
@@
3.3 DocT5Query vs. Llama 2 Expansions
In this section, we compare the quality of document expansions generated by DocT5Query and Llama 2 using BM25 scores on the MS MARCO Dev Queries dataset. The effectiveness metrics are summarized in Table 3 for varying numbers of queries appended to each document in the corpus.
@@
Table 3: Comparison of effectiveness metrics (MRR@10 and Recall@1000) between DocT5Query and Llama 2 expansions for varying numbers of appended queries using BM25 scores on MS MARCO Dev Queries.
MRR@10 Recall@1000
n DocT5Query Llama 2 DocT5Query Llama 2
5 25.9 25.4 92.9 94.0
10 26.5 26.5 93.9 94.8
20 27.2 25.4 94.4 95.4
40 27.7 27.4 94.7 95.5
80 27.8 27.4 94.5 95.6
3.4 Recall at Different Depths
Table 4 presents the recall metrics at various depths on the MS MARCO Dev Queries dataset for four different models: uniCOIL (with TILDE document expansion), SPLADE (which refers to CoCondenser-EnsembleDistil in [23]), DeepImpact  and DeeperImpact. We notice that SPLADE consistently achieves the highest recall at all the depths. However, DeeperImpact—which offers much faster query processing—also demonstrates strong effectiveness, keeping ahead of uniCOIL in all the depths.
Remarkably, DeeperImpact attains a recall at depth 400 that is equivalent to the Recall@1000 achieved by the original DeepImpact model. This highlights the effectiveness of DeeperImpact in retrieving relevant documents at shallower depths, demonstrating its efficiency in reducing the number of retrieved documents necessary to achieve high recall.
Table 4: Comparative analysis of recall at various depths on MS MARCO Dev Queries for uniCOIL, SPLADE, DeepImpact, and DeeperImpact.